ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

17. Control Theory and Optimization

1991 MS Classification: 49-04, 65C20, 93C95 Pesch, Hans Josef; Hinsberger, Henrik, Clausthal University of Technology, Germany Optimal Control of Large Scale DAE Systems in Chemical Engineering In chemical process control challenging problems for numerical optimal control methods arise very often. Those problems are generally governed by large scale differential-algebraic equations. A new direct multiple shooting method for the numerical solution of those problems is presented. The method exploits the particular structure of the system equations and allows to formulate the control problem as a multistage process so that state and/or control space dimension as well as objective function and path constraints may vary from stage to stage. Moreover, this is advantageous for parallelization techniques, too, and leads to a new divide and conquer method. Numerical results are presented for some real life problems.